# Let’s do the time-warp-attend: Learning topological invariants of dynamical systems

### 19 Jan 2024 v2

@@ 4.2 Recovering bifurcation boundaries
Figure 2: Inference of bifurcation boundary. Each system is plotted at coordinates of its simulated parameter values and colored according to its average cycle prediction across 50 training re-runs. The true parametric bifurcation boundary is plotted in black. Colors are scaled between a perfect cycle classification (red) and perfect point classification (blue) with white at evenly split classification.
@@ 4.3 Predicting oscillatory regulation of the repressilator system
In order to demonstrate the viability of our method for higher-dimensional systems under realistic noise conditions, we apply our topological approach to a computational model in a well-studied biological setting, the so-called repressilator model (Elowitz & Leibler, 2000). The repressilator system is a synthetic, genetic regulatory network that was introduced by Elowitz and Liebler into bacteria to induce oscillatory regulation of gene expression Fig. 3. The system is modeled by the mRNA and protein counts of three genes (TetR, LacI and $\lambda$ phage cI) that inhibit each other in a cyclic manner (see Appendix 7.7 for governing equations). Theoretically and experimentally, the existence and characteristics of oscillations have been shown to depend on numerous factors like the rates of transcription, translation, and degradation, as well as intrinsic and extrinsic noise (Verdugo, 2018; Potvin-Trottier et al., 2016). Here we test our method in classifying a regime of the repressilator which has been shown to be undergoing a Hopf bifurcation, though, unlike the previous complex simulated systems, it spans six dimensions and is constructed from noisy cellular time-series.
To generate 2D vector field data, we consider trajectories of cellular gene expression measured by protein levels of TetR and LacI (as in Fig. 3, middle). Different instances of simulations share the initial condition but vary by the rate of transcription, $\alpha$, and by the ratio of protein and mRNA degradation rates, $\beta$. We sample 100 cell states along the temporal trajectory and add Gaussian noise of $\sigma=0.5$ to the gene expression states (to protein and mRNA levels) to model intrinsic noise. For each set of measurements, we interpolate an input 64x64 vector field, based on their velocities across TetR-LacI protein phase space. For further details of the data generation process see Appendix 7.7. We predict for each vector field its dynamical class, achieving 87% classification accuracy and proximate bifurcation boundary to the analytically-derived boundary (Fig. 3, right (Verdugo, 2018)). In Fig. A13 we show prototypes of trajectories across the $\alpha-\beta$ parameter space, where we see that where we fail to classify a point attractor, trajectories of cells loop closely in the TetR-LacI protein plane and generate data samples that highly resemble trajectories with a periodic attractor rather than with a point attractor.
Figure 3: Oscillations in repressilator system. Example of oscillations in the repressilator system viewed as a time-series (left) or projected onto TetR-LacI protein plain (middle). (right) Predicted oscillatory behavior by our model output upon varying transcription rate without repression ($\alpha$) and the ratio of protein and mRNA degradation rates ($\beta$). The theoretical boundary is marked in black. See details in Appendix 7.7.
@@ 4.4 Distinguishing dynamical regimes in single-cell data of the pancreas
We focus on the process of pancreas endocrinogenesis where Ductal and Ngn3+ precursor cells rapidly proliferate to give rise to functional cell types Alpha, Beta, Delta and Epsilon (Bastidas-Ponce et al., 2019). We examine a single-cell RNA-sequencing dataset where the expression of over 2,000 genes are measured per cell, for a total of 3,600 cells. While the RNA expression profile, providing a gene expression vector per cell, provides for each cell its location within gene expression space, RNA velocity (Bergen et al., 2021; Gorin et al., 2022) can be used to compute the effective velocity of each cell (based on differences in spliced and unspliced RNA counts). We use scVelo (Bergen et al., 2020) to infer cellular velocities in the high-dimensional gene expression space and to project them to a 2D UMAP representation (Fig. 4, left). Simultaneously, we also compute for each cell a ‘cell-cycle score’, which captures the extent of expression of cell-cycle genes (in phase S and in grouped G2 and M phases, Fig. 4, middle; see Appendix 7.8, Bergen et al. (2020); Tirosh et al. (2016)) . For each set of cells, we interpolate an input 64x64 vector field, based on their RNA velocities, over the 2D UMAP representation.
To generate multiple frames from the trajectory, we randomly split cells of each cell type into groups of at least 50 cells resulting in 69 (sparse, noisy) vector fields with corresponding cell-cycle score (Fig. 4, middle, see details in Appendix 7.8). We then used our model (still trained solely on simulated augmented simple oscillator data) to characterize the dynamic regimes of the cellular behavior. Our model can successfully identify proliferative, cyclic behavior within the single-cell data, as reflected by 94% accuracy in predicting a cyclic behavior matching the high cell cycle score (Fig. 4, right), exceeding the prediction accuracy reached by alternative baselines (Critical Points: 54%, Lyapunov: 49%, Phase2vec: 30%, Autoencoder: 30%). As opposed to the cell-cycle score, our model does not use any prior knowledge regarding cell-cycle related genes, and is only based on the structural features of the cellular vector fields, and therefore can also be employed when such prior knowledge is not available.
Figure 4: Proliferation-to-differentiation transition in pancreas development. (left) Gene expression velocity of cells undergoing pancreas endocrinogenesis in UMAP space. Colors distinguish between different cell types. (middle) Cell-cycle score based on the fraction of cells with high S-phase gene expression (with S-score $>$ 0.4; see Appendix 7.8). (right) Predicted cell-cycle score computed as the average classification prediction across 50 training re-runs.
@@ 5 Limitations
Our framework learns to be invariant to geometric (as opposed to topological) properties of systems through topological data augmentation and convolutional attention, achieving competitive results in bifurcation detection across real-world systems. Nevertheless, we note here several important limitations of our method in its current form. The method is currently confined to the important but circumscribed setting of two-dimensional dynamics. Scaling up to much higher dimensions in these contexts presents technical challenges like the exponential growth of input size with dimension, as well as theoretical challenges such as the lack of structural stability in higher dimensions and the introduction of chaotic regimes. Since inherently low-dimensional dynamics often underlie high-dimensional measurements, especially in biology (Xiong & Garfinkel, 2023; Khona & Fiete, 2022), high dimensions can be mitigated by hardwired, or optimized dimensionality reduction methods as we deploy for the six-dimensional repressilator and for analyzing over 2,000 gene expression dynamics of the pancreas single-cell dataset. Additionally, we used a supervised, multi-class setup so that the classifier can predict that both a limit cycle and a fixed point co-exist. Empirically, we find that our method learns that the point and periodic attractors are exclusive, in fact, the learned logits consistently negate each other. For emphasis, we tested our trained model on samples representing a subcritical Hopf bifurcation having a regime with point and periodic attractors whose logits were of intermediate value but still negations, see Appendix 7.6. In the future, we aim to extend to combinatorial annotations of invariant sets, as elaborated in the discussion.
@@ 6 Discussion
Our work suggests numerous next steps including further translating our prior knowledge of topological properties, like the importance of critical points, into either explicit features or trainable, physics-informed network components (Karniadakis et al., 2021). Another goal is to broaden the scope to encompass a wider range of invariance types, combinatorial comparisons of invariant sets (e.g., via approximation of Conley-Morse graphs), and extensions to stronger notions of equivalence, such as topological conjugacy. To accommodate more versatile invariances and bifurcations, future work could extend our framework to the unsupervised case. For example, given that our data augmentations produce samples which are in the same dynamical equivalence class, a contrastive loss could be used to shape the feature space so that bifurcation classes are represented implicitly during training. Or, while challenging, following Skufca & Bollt (2008), a purely unsupervised approach could be used to explicitly construct warpings between systems in order to characterize equivalence classes.
Our work demonstrates the potential for scaling up to applications such as identifying dynamical regimes and underlying driving factors across both biological and physical applications.

### 21 Mar 2024 v3

@@ 4.2 Recovering bifurcation boundaries
Figure 2: Inference of bifurcation boundaries. Each system is plotted at coordinates of its simulated parameter values and colored according to its average cycle prediction across 50 training re-runs. The true parametric bifurcation boundary is plotted in black. Colors are scaled between a perfect cycle classification (red) and perfect point classification (blue) with white at evenly split classification.
@@ 4.3 Predicting oscillatory regulation of the repressilator system
In order to demonstrate the viability of our method for higher-dimensional systems under realistic noise conditions, we apply our topological approach to a computational model in a well-studied biological setting, the so-called repressilator model (Elowitz & Leibler, 2000). The repressilator system is a synthetic, genetic regulatory network that was introduced by Elowitz and Liebler into bacteria to induce oscillatory regulation of gene expression (Fig. 3). The system is modeled by the mRNA and protein counts of three genes (TetR, LacI and $\lambda$ phage cI) that inhibit each other in a cyclic manner (see Appendix 7.8 for governing equations). Theoretically and experimentally, the existence and characteristics of oscillations have been shown to depend on numerous factors like the rates of transcription, translation, and degradation, as well as intrinsic and extrinsic noise (Verdugo, 2018; Potvin-Trottier et al., 2016). Here we test our method in classifying a regime of the repressilator which has been shown to undergo a Hopf bifurcation, though, unlike the previous complex simulated systems, it spans six dimensions and is constructed from noisy cellular time-series.
To generate 2D vector field data, we consider trajectories of cellular gene expression measured by protein levels of TetR and LacI (Fig. 3, middle). Different instances of simulations share the initial condition but vary by the rate of transcription, $\alpha$, and by the ratio of protein and mRNA degradation rates, $\beta$. We sample 100 cell states along the temporal trajectory and add Gaussian noise of $\sigma=0.5$ to the gene expression states (to protein and mRNA levels). For each set of measurements, we interpolate an input 64x64 vector field, based on their velocities across TetR-LacI protein phase space, see examples at Fig. A14 and Appendix 7.8 for further details. We predict for each vector field its dynamical class, achieving 87% classification accuracy and proximate bifurcation boundary to the analytically-derived boundary (Fig. 3, right (Verdugo, 2018)). In Fig. A13 we show prototypes of trajectories across the $\alpha-\beta$ parameter space, where we see that where we fail to classify a point attractor, trajectories of cells loop closely in the TetR-LacI protein plane and generate data samples that highly resemble trajectories with a periodic attractor rather than with a point attractor.
Figure 3: Oscillations in a repressilator system. Example of oscillations in the repressilator system viewed as a time-series (left) or projected onto TetR-LacI protein plain (middle). (right) Predicted oscillatory behavior by our model output upon varying transcription rate without repression ($\alpha$) and the ratio of protein and mRNA degradation rates ($\beta$). The theoretical boundary is marked in black. See details in Appendix 7.8.
@@ 4.4 Distinguishing dynamical regimes in single-cell data of the pancreas
We focus on the process of pancreas endocrinogenesis where Ductal and Ngn3+ precursor cells rapidly proliferate to give rise to functional cell types Alpha, Beta, Delta and Epsilon (Bastidas-Ponce et al., 2019); In that study, a single-cell RNA-sequencing dataset was collected where the expression of over 2,000 genes are measured per cell, for a total of 3,600 cells (Fig. 4, left). While the RNA expression profile, providing a gene expression vector per cell, provides for each cell its location within gene expression space, RNA velocity (Bergen et al., 2021; Gorin et al., 2022) can be used to compute the effective velocity of each cell (based on differences in spliced and unspliced RNA counts). We use scVelo (Bergen et al., 2020) to infer cellular velocities in the high-dimensional gene expression space and to project them to a 2D UMAP representation. Simultaneously, we also compute for each cell a ‘cell-cycle score’, which captures the extent of expression of cell-cycle genes (in phase S and in grouped G2 and M phases, Fig. 4, middle; see Appendix 7.9, Bergen et al. (2020); Tirosh et al. (2016)) . For each set of cells, we interpolate an input 64x64 vector field, based on their RNA velocities, over the 2D UMAP representation.
To generate multiple frames from the trajectory, we randomly split cells of each cell type into groups of at least 50 cells resulting in 69 (sparse, noisy) vector fields with corresponding cell-cycle score (Fig. 4, middle, Appendix 7.9). We then used our model (still trained solely on simulated augmented simple oscillator data) to characterize the dynamic regimes of the cellular behavior. Our model can successfully identify proliferative, cyclic behavior within the single-cell data, as reflected by 94% accuracy in predicting a cyclic behavior matching high cell cycle scores (Fig. 4, right), substantially exceeding the prediction accuracy reached by alternative baselines (Critical Points: 54%, Lyapunov: 49%, Phase2vec: 30%, Autoencoder: 30%). As opposed to the cell-cycle score, our model does not use any prior knowledge regarding cell-cycle related genes, and is only based on the structural features of the cellular vector fields, and therefore can also be employed when such prior knowledge is unavailable.
Figure 4: Proliferation-to-differentiation transition in pancreas development. (left) Gene expression velocity of cells undergoing pancreas endocrinogenesis in UMAP space. Colors distinguish between different cell types. (middle) Cell-cycle score based on the fraction of cells with high S-phase gene expression (with S-score $>$ 0.4). (right) Predicted cell-cycle score computed as the average classification prediction across 50 training re-runs. See details in Appendix 7.9.
@@ 5 Limitations
Our framework learns to be invariant to geometric (as opposed to topological) properties of systems through topological data augmentation and convolutional attention, achieving competitive results in bifurcation detection across real-world systems. Nevertheless, we note here several important limitations of our method in its current form. The method is currently confined to the important but circumscribed setting of two-dimensional dynamics. Scaling up to much higher dimensions in these contexts presents technical challenges like the exponential growth of input size with dimension, as well as theoretical challenges such as the lack of structural stability in higher dimensions and the introduction of chaotic regimes. Since inherently low-dimensional dynamics often underlie high-dimensional measurements, especially in biology (Xiong & Garfinkel, 2023; Khona & Fiete, 2022), high dimensions can be mitigated by hardwired, or optimized dimensionality reduction methods as we deploy for the six-dimensional repressilator and for analyzing over 2,000 gene expression (dimensional) dynamics of the pancreas single-cell dataset. Additionally, we used a supervised, multi-class setup so that the classifier can predict that both a limit cycle and a fixed point co-exist. Empirically, we find that our method learns that the point and periodic attractors are exclusive, in fact, the learned logits consistently negate each other. For emphasis, we tested our trained model on samples representing a subcritical Hopf bifurcation having a regime with point and periodic attractors whose logits were of intermediate value but still negations, see Appendix 7.7. In the future, we aim to extend to combinatorial annotations of invariant sets, as elaborated in the discussion.
@@ 6 Discussion
Our work suggests numerous next steps including further translating our prior knowledge of topological properties, like the importance of critical points, into either explicit features or trainable, physics-informed network components (Karniadakis et al., 2021). Another goal is to broaden the scope to encompass a wider range of invariance types, combinatorial comparisons of invariant sets (e.g., via approximation of Conley-Morse graphs), and extensions to stronger notions of equivalence, such as topological conjugacy. To accommodate more versatile invariances and bifurcations, future work could extend our framework to the unsupervised case. For example, given that our data augmentations produce samples which are in the same dynamical equivalence class, a contrastive loss could be used to shape the feature space so that dynamical classes are represented implicitly during training. Or, while challenging, following Skufca & Bollt (2008), a purely unsupervised approach could be used to explicitly construct warpings between systems in order to characterize equivalence classes. Our work demonstrates the potential for scaling up to applications such as identifying dynamical regimes and underlying driving factors across both biological and physical applications.